How many ways can seven people be seated in a row if two of them refuse to sit next to each other?
Answer by stanbon(75887)

How many ways can seven people be seated in a row if two of them refuse to sit next to each other?
------------------------
Place those two people:
7 ways to place the 1st person
5 wyas to place the 2nd person
Place the remaining 5 people - 5! ways
-----------------------
Total ways = 7*5*5! = 35*120 = 4200 ways
=====================
Cheers,
Stan H.